OpenStudy (anonymous):

Solve for x lnx +ln(x-1) = 1

OpenStudy (anonymous):

value of log of a base a is 1 so we can write 1 as ln e equation becomes x(x-1)=e solve

OpenStudy (anonymous):

hmm, honestly I am not sure of where to go from there, I was thinking it was x^2-1 = e and then x^2 = e+1 and then x = sqrt(e+1) but I forgot to multiple the x by the 1 too so it would have been x^2 - x = e. Not dure where to go from here or if I should have multiplied the 2.

OpenStudy (anonymous):

Not sure*

OpenStudy (anonymous):

use e as a constant value..u can't skip x like dat. u hav to consider it.. make it x^2-x-e=0 and solve the quadratic equation

OpenStudy (anonymous):

When I plug in e, should I plug in the value 2.7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay I will try it out

OpenStudy (anonymous):

Okay I got x = 2.22287 and x = -1.22287 , how can I check if the answers are right?

OpenStudy (anonymous):

x cannot be -ve so x= 2.22 is ur ans

OpenStudy (anonymous):

I answered it in my online thing and it said it was wrong :/

OpenStudy (anonymous):

dats weird...can u post the image of the question?

OpenStudy (anonymous):

I am trying to post it but it is not working, can you try to work out the question to see if you get the same answer?

OpenStudy (anonymous):

even i am getting 2.22

OpenStudy (anonymous):

Are you sure the method is correct?

OpenStudy (anonymous):

Here I will make a new post